OpenStudy (hba):

if 2x^-5=64,what is the value of x ?

OpenStudy (hba):

Please check out the Coc- http://openstudy.com/code-of-conduct,You are not allowed to give the answers directly. -_-

OpenStudy (hba):

@pathosdebater Please explain ?

OpenStudy (anonymous):

2x^-5=64 x^-5=32 1/(x^5)=2^5 x^5=(1/2)^5 x=1/2

OpenStudy (hba):

So this is the real root :)

OpenStudy (anonymous):

yes

OpenStudy (hba):

I think you guys forgot complex roots,anyways thanks a lot.

OpenStudy (anonymous):

You want the complex roots?

OpenStudy (hba):

No thanks.

OpenStudy (hba):

I know how to do it.